OpenStudy (mgourab):

Hello, I am unable to understand a question, can any1 please explain me the concept and some tricks to quicky understand the sam? : " if sum of 3 no's is 65. the ratio of 1st & 2nd is (5:3) and ratio of 2nd & 3rd is (7:3), then what is the second no?

OpenStudy (wolf1728):

You need 3 equations to solve this: F + S + T = 65 F/S = 5/3 S/T = 7/3 Solve for the 3 unknowns and you'll have all the values for the First Second and Third number.

OpenStudy (unklerhaukus):

Solve the second equation for F, Sole the third equation for T, Plug these into the first equation, such that you have an equation with only S. Solve for S.

OpenStudy (proheiper):

^

OpenStudy (arnavguddu):

A:B = 5:3 = 5/3 : 1 B:C = 7:3 = 1 : 3/7 A:B:C = 5/3 : 1 : 3/7 let => A = (5/3)* k B = k C = (3/7) * k A + B + C = (5/3) * k + k + (3/7) * k = 65 now find k now with k known, A, B, C can be found easily